40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Dec 18 16:21:08 2014, ericq, Update, Computer Scripts / Programs, scripts not being backed up! 
    Reply  Fri Dec 19 16:59:32 2014, ericq, Update, Computer Scripts / Programs, FSS Slow servo moved to megatron 
       Reply  Fri Dec 19 20:44:23 2014, Jenne, Update, Computer Scripts / Programs, FSS Slow servo moved to megatron 
          Reply  Sat Dec 20 00:00:03 2014, ericq, Update, Computer Scripts / Programs, FSS Slow servo moved to megatron 
             Reply  Tue Dec 23 18:43:33 2014, diego, Update, Computer Scripts / Programs, FSS Slow servo moved to megatron 
                Reply  Fri Dec 26 18:20:42 2014, rana, Update, Computer Scripts / Programs, FSS Slow servo thresh change 
Message ID: 10844     Entry time: Fri Dec 26 18:20:42 2014     In reply to: 10840
Author: rana 
Type: Update 
Category: Computer Scripts / Programs 
Subject: FSS Slow servo thresh change 

Quote:

 

In the plot it is shown the behaviour of the PSL-FSS_SLOWDC signal during the last week; the blue rectangle marks an approximate estimate of the time when the scripts were moved to megatron. Apart from the bad things that happened on Friday during the big crash, and the work ongoing since yesterday, it seems that something is not working well. The scripts on megatron are actually running, but I'll try and have a look at it.

I guessed that what was happening was that the SLOW servo settings were not restored to the right values after the code movements / reboots. The ON threshold for the servo was set at +6 counts and the channel is MC TRANS. Since the ADC noise on that channel is ~50 counts, this means that the servo keeps pushing the laser temperature off in some direction when the MC is unlocked.

I reset the threshold to +6666 counts (the aligned MC transmission is ~16000 for the  TEM00 mode) so that it only turns on when we're in a good locked state.

ELOG V3.1.3-